OpenStudy (lilai3):

How do you write: 42 decreased by twice a number in algebraic expression WILL REWARD MEDAL

OpenStudy (raden):

42 - 2x (in x)

OpenStudy (lilai3):

please explain

OpenStudy (lilai3):

why isn't it 2x-42?

OpenStudy (raden):

that's just a translater from words to expression algebraic

OpenStudy (lilai3):

oh. what if it says... the sum of the cube of a number and 12.... is it n^3 +12?

OpenStudy (raden):

if it says : 42 decreased "from" twice a number, yeah it is 2x - 4 the sum of the cube of a number and 12, yeah you are right it is n^3 + 12

OpenStudy (raden):

* 2x - 42 i mean

OpenStudy (lilai3):

oh. thank you very much.

OpenStudy (raden):

you're welcome

OpenStudy (lilai3):

also, i have one more problem

OpenStudy (lilai3):

name the property used in each equation. then find the value of n. 7y+y=7y+ny

OpenStudy (raden):

im not sure is the answer " reflexive or symmetric property" im forget this, sorry :)

OpenStudy (raden):

@rizwan_uet , you know this ?

OpenStudy (lilai3):

oh.at least you told me one of the,. thanks. i thought it was commutitive.. LOL. uhm... what is n?

OpenStudy (anonymous):

isnt it obeys distributive property??? @RadEn

OpenStudy (raden):

a+b = b+a (it is commutative) a(b+c) = ab + ac (it is distributive property)

OpenStudy (lilai3):

so is it reflexive, since symmetry it has to be 7y+y=ny+7y, but it's not... so is it reflexive?

OpenStudy (anonymous):

if we take y common on both sides we get y(7+1) =y (7+n) now using distributive property on each side we get 7y+y=7y+ny

OpenStudy (lilai3):

i don't get it....

OpenStudy (raden):

me too, haha...

OpenStudy (anonymous):

you have this equation 7y+y=7y+ny now some property of real number was applied to a certain equation to get 7y+y=7y+ny my guess was that the equation was in this form y(7+1) =y (7+n) now applying distributive property in both sides we get 7y+y=7y+ny sorry for my weak english LOL

OpenStudy (lilai3):

lol. it wasn't weak. :P ...but then what is n? because you have to find the value of that.... is it 1? or not... idunno

OpenStudy (anonymous):

7y+y=7y+ny the only value that n can hold is 1 to get both sides similer of the above equation

OpenStudy (lilai3):

so i'm right? it's distributive and 1?

OpenStudy (anonymous):

for the value of n you are absolutly right but for distributive property it was my state of arguments

OpenStudy (lilai3):

oh. okay... can you also do another problem...let's say(random): (6+n)x+15x

OpenStudy (anonymous):

(6+n)x+15x using multiplicative property property we get x(6+n)+15 now using distributive property 6x+nx +15x

OpenStudy (lilai3):

why did you add 15 x to the equation?

OpenStudy (anonymous):

its the part of equation

OpenStudy (lilai3):

then where's the equal sign?

OpenStudy (anonymous):

hmm wait

OpenStudy (lilai3):

okay

OpenStudy (anonymous):

(6+n)x+15x using multiplicative property property we get = x(6+n)+15x now using distributive property = 6x+nx +15x

OpenStudy (anonymous):

does this make sense now??

OpenStudy (lilai3):

yes, so it is distributive property, and the n is 9, right?

OpenStudy (anonymous):

basically you wrote this at first (6+n)x+15x now its not an equation until you mention some equality sign here now you cant find n in the absence of equality sign

OpenStudy (anonymous):

sorry i lost internet connection

OpenStudy (lilai3):

it's alright, and thank you!

OpenStudy (anonymous):

welcome